OpenStudy (anonymous):

dwyane is taking a drawing class. the drawing pencils cost $.97 a piece and a sketchbook cost $5.95. he spent a total of of $11.77. write and solve a equation to find the number "n" of pencils he bought.

OpenStudy (anonymous):

I got n= 6

OpenStudy (anonymous):

if ur typing i forgot im not getting your sollution

OpenStudy (anonymous):

hello r u there

OpenStudy (anonymous):

well you have total cost x = pencil = ( want to see how many you need) y = drawing board = (only have one) k = constant ( result of what you have, value at the end you work with) x = .97 y = 5.95 k = 11.77 .97 x + 5.95 y = 11. 77 [ .97 x + 5.95 y = 11. 77 ] 100 97 x + 595 y = 11 77 97 x + 595 + (- 595 ) = 1177 + (-595) [ 97x = 582 ] 1/ 97 x = 6 (i multiplied by 100 to get rid of decimals) ya its 6 pencils

OpenStudy (anonymous):

there was a lot to type sorry

OpenStudy (anonymous):

its

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it ok what i put up there?

OpenStudy (anonymous):

yes thanks a hole lot for your timw i understand !!!